OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x - 2y = -4 2x - y = 1

OpenStudy (anonymous):

solve this the same way as the last question you asked

OpenStudy (anonymous):

eed help

OpenStudy (anonymous):

did you understand the answer to your last question?

OpenStudy (anonymous):

y=2x-1 therefore x-4x+2 = -4 -3x = -6 therefore x=2 thus y=3 (2,3)

jhonyy9 (jhonyy9):

x+2y=7 / 2 2x+y=8 2x+4y=14 2x+y=8 --------- -- 3y=6 y=6/3 =2 x+2y=7 x+4=7 x=7-4=3 x=3 --- solutions are (3,2)